OpenStudy (anonymous):

can you help me simplify this? I dunno why I'm having troubles: (3a - 6 - 3b+6 /ab-2a-2b+4)/b-a

OpenStudy (whpalmer4):

Is that supposed to be \[\frac{(\frac{3a - 6 -3b + 6 } {ab - 2a - 2b +4 } )} {b-a }\]

OpenStudy (anonymous):

yeah

OpenStudy (whpalmer4):

Well, I would factor everything, and then remember that dividing a fraction by something is the same as multiplying the fraction by 1/something, so that you'd have \[\frac{3a-6-3b+6}{a b - 2a -2b + 4}*\frac{1}{b-a}\]and then you'd be able to do some cancelation to simplify

OpenStudy (anonymous):

so i got it down to -3(-a+b) divided by (a-2)(b-2) multiplied by 1/b-a. the top -a+b can be crossed out by 1/b-a but I don't know what more I can do

OpenStudy (anonymous):

am i doing something totally wrong?

OpenStudy (whpalmer4):

Do you expect the end result will be something very simple? \[-\frac{3}{(a-2)(b-2)}\] is as good as you can do, I believe.

OpenStudy (jhannybean):

Do you understand @whpalmer4 's method?

OpenStudy (whpalmer4):

the fraction becomes \[\frac{3a -6 - 3b + 6}{ab - 2a - ab + 4}*\frac{1}{b-a} = \frac{3(a-b)}{(a b - 2a -2b + 4)(b-a)} \]\[= \frac{-3\cancel{(b-a)}}{\cancel{(b-a)}(ab - 2a - 2b + 4) } = -\frac{3}{(a-2)(b-2)}\]

OpenStudy (jhannybean):

Yep.to flip a-b multiply the terms by -1. easier to cancel.

OpenStudy (anonymous):

thanks guys!